# DecompDreamer: Advancing Structured 3D Asset Generation with Multi-Object Decomposition and Gaussian Splatting

### v1

@@ Abstract
Text-to-3D generation saw dramatic advances in recent years by leveraging Text-to-Image models. However, most existing techniques struggle with compositional prompts, which describe multiple objects and their spatial relationships. They often fail to capture fine-grained inter-object interactions. We introduce DecompDreamer, a Gaussian splatting-based training routine designed to generate high-quality 3D compositions from such complex prompts. DecompDreamer leverages Vision-Language Models (VLMs) to decompose scenes into structured components and their relationships. We propose a progressive optimization strategy that first prioritizes joint relationship modeling before gradually shifting toward targeted object refinement. Our qualitative and quantitative evaluations against state-of-the-art text-to-3D models demonstrate that DecompDreamer effectively generates intricate 3D compositions with superior object disentanglement, offering enhanced control and flexibility in 3D generation. Project page : https://decompdreamer3d.github.io
Figure 1: Illustration of high-quality, complex 3D scenes generated by DecompDreamer for prompts containing 6-11 objects. Existing methods struggle to generate all objects accurately and fail to maintain correct spatial relationships.
footnotetext: * Equal Contribution.footnotetext: $\dagger$ Corresponding author: unath@asu.edu
@@ 1 Introduction
Figure 2: (a) Illustration of feature-object mismatch. The image generated by StableDiffusion 2.1 using the prompt ”A blue bird sitting on a black branch of a white tree” (b) 3D representations generated by LucidDreamer [20] and DecompDreamer.
3D representations play a crucial role in many applications including perception and autonomy, as well as in the media industries, enabling the creation of immersive and realistic digital environments. These industries strive to replicate real-world complexity by incorporating high-fidelity 3D objects that interact dynamically within virtual spaces. Real-world scenes consist of objects and agents with complex spatial relationships and potential interactions, with their surroundings, that can only be described by intricate semantic relationships. Recent methods employ 2D lifting techniques [6, 21, 24, 32, 40] to generate 3D models from text using pre-trained diffusion models [34, 29, 35] and Score Distillation Sampling (SDS) [32]. While effective for single objects with simple text prompts, these methods struggle with compositional prompts involving multiple objects, with quality significantly degrading as inter-object relationships increase. This decline is primarily due to (1) feature-object mismatch and (2) poor modeling of 3D spatial relationships.
Feature-object mismatch arises as a fundamental limitation of text-image embeddings [33]. As text complexity increases, matching attributes to their corresponding objects becomes increasingly challenging. For instance, in the prompt “A blue bird sitting on a black branch of a white tree” (Figure 2(a)), the text-to-image model confuses the attributes “blue” and “white,” resulting in a bird with blue-white coloration instead of the intended output. Furthermore, as the number of inter-object relationships grows, existing methods fail to maintain spatial coherence, leading to incorrect placements, overlapping artifacts, or missing contextual dependencies. For example, in Figure 2(b), the generated 3D representation by LucidDreamer [20] deteriorates as the complexity and number of inter-object relationships increase.
While generating a high-fidelity 3D representation corresponding to compositional text prompts is crucial, it is equally important to break the scene down into smaller, well-defined components. A monolithic 3D output limits flexibility, making it difficult to refine individual objects or modify their attributes without affecting the entire scene. A decomposed representation provides greater control, allowing artists and developers to tweak specific parts while preserving overall scene coherence.
In this paper, our goal is to generate decomposable 3D representations corresponding to complex text prompts involving intricate inter-object relationships. Let us consider how an artist might approach this problem [9]. An expert 3D modeler tasked with designing a 3D scene consisting of multiple objects would likely begin by identifying the number of simple objects, estimating their approximate relative positions, and then creating meshes for each object while refining their spatial relationships. Following this simple yet effective strategy, we propose DecompDreamer, a Gaussian-splatting-based optimization routine that decomposes complex 3D representations into smaller components and progressively refines them. Our approach is built on three core design choices:
Structured 3D Decomposition and Initialization: Given a text prompt, we use a Vision-Language Model (VLM) to decompose the scene into individual objects, infer their spatial arrangement, and establish inter-object relationships. Each object is represented using 3D Gaussians, enabling structured parameter initialization via coarse 3D point clouds generated by Point-E [28]. Unlike implicit representations such as NeRF [25], which lacks object-level control, this approach ensures a spatially consistent and well-structured starting point.
Progressive Optimization: Existing SDS-based methods optimize entire scenes at once [6, 21, 24, 32, 40, 43] or iteratively refine objects, relationships and full scene [13, 14]; often leading to entangled representations and inconsistencies. While effective for simple objects or those with few inter-object relationships, their performance degrades significantly as the number of relationships increases. In this paper, we propose a progressive optimization strategy inspired by artistic workflows. Unlike existing methods, we do not primarily focus on optimizing the entire scene at once. Instead, we emphasize modeling inter-object relationships and refining individual objects, with occasional scene-level optimization. We start with joint optimization of objects with respect to their relationships, then gradually shift focus toward individual object refinement. Optimizing objects in a compositional environment is significantly more challenging than in isolation as it requires accounting for spatial dependencies and relative viewing direction. To address this, we incorporate object-view-aware optimization, improving object disentanglement and ensuring a coherent 3D structure. Once the scene is well-formed, we apply targeted optimization within the relational context to refine fine-grained object details while preserving spatial coherence.
Flow Matching Distillation: To effectively model objects and their relationships, we extend diffusion-based SDS to flow-based SDS, leveraging rectified flow models [1, 2] for smoother and more controlled optimization. Flow matching enables stable and efficient optimization by learning a direct vector field, eliminating stochastic noise and improving convergence.
With these advancements, DecompDreamer is capable of generating high-quality compositional 3D objects (Figure 1, 4) while achieving enhanced disentanglement for individual objects (Figure 5). We evaluate DecompDreamer through both qualitative and quantitative analyses. A user study indicates that DecompDreamer outperforms existing models, including Gala3D [43], GraphDreamer [13], LucidDreamer [20], MVDream [36], and HiFA [45], in terms of image quality and semantic alignment. Additionally, we perform quantitative evaluations using CLIP score [33], and Pick-A-Pic [19] on 30 compositional text prompts containing 2 to 11 objects, demonstrating DecompDreamer’s superior performance in compositional 3D generation.
@@
  1. 1.
    We introduce DecompDreamer, a 3D Gaussian-based training routine that decomposes complex 3D representations into smaller, manageable parts and applies progressive optimization to refine relationships and individual objects, using a novel flow matching distillation approach.
  2. 2.
    DecompDreamer achieves strong disentanglement across individual object components using view-aware optimization, enhancing interpretability and controllability.
  3. 3.
    We extensively validate our approach using a diverse set of complex text prompts and benchmark it both qualitatively and quantitatively against several state-of-the-art 3D generators [21, 6, 40, 13, 43]. Empirical evaluations demonstrate that DecompDreamer consistently produces well-structured, high-fidelity compositional 3D objects.
@@ 2 Related Work
Text-to-3D Generation converts textual descriptions into 3D objects and is broadly categorized into 3D supervised and 2D lifting techniques. 3D supervised methods  [8, 17, 28, 41] use paired text and 3D data for training, enabling high-quality 3D generation. However, they struggle to generalize to compositional scenes, as most 3D datasets  [10, 23] focus on single-object models. In contrast, 2D lifting techniques  [32, 21, 6, 40, 44, 16, 7] refine 3D representations using pre-trained 2D diffusion models. DreamFusion  [32] introduced Score Distillation Sampling (SDS) to align 3D content with a 2D diffusion prior, inspiring further advancements  [20, 21, 39, 38, 44]. However, these methods struggle with compositional 3D generation as they do not model inter-object relationships individually but instead treat the entire scene as a single entity, leading to poor disentanglement of objects and relationships.
To address this problem, recent methods [11, 4, 31, 42] incorporate 3D layouts with non-overlapping bounding boxes. While effective for structured scenes, they struggle with overlapping objects. GraphDreamer [13] addresses this issue by using a scene graph to represent 3D scenes, separately modeling objects and their relationships. However, it relies on position embeddings to map 3D locations to objects, making optimization challenging. Consequently, it struggles to scale to scenes with a large number of objects and suffers from poor disentanglement. Gala3D [43] enables compositional 3D generation by optimizing Gaussians for individual objects and the overall scene. However, it struggles with complex interactions, particularly overlapping objects, leading to spatial inaccuracies. CompGS [14] enhances initialization using pre-trained text-to-3D models yet struggles with intricate relationships. We propose DecompDreamer to address these challenges and enable the generation of compositional scenes with a large number of objects and intricate relationships. Similar to Gala3D and CompGS, we use Gaussians as the 3D representation. However, instead of jointly optimizing objects and the entire scene, we adopt a progressive optimization approach, transitioning from inter-object relationships to individual object refinement for more structured and disentangled 3D generation.
Rectified Flow Models (RFM)  [3, 22] have recently gained attention as an alternative generative approach to diffusion models. Unlike traditional diffusion models  [15, 37], which rely on stochastic sampling via a stepwise denoising process, RFMs directly learn smooth, continuous transport between data distributions, eliminating iterative noise perturbations and leading to more efficient sampling and improved generation quality. Recent works  [12] have demonstrated the effectiveness of RFMs for large-scale image and video generation. In this paper, we apply RFMs to 3D generation tasks, showcasing their capabilities.
@@ 3 Preliminaries
Gaussian Splatting  [18] models a scene with a collection of 3D Gaussians, characterized by means, covariances, colors, and opacities. These Gaussians are projected onto the image plane as 2D Gaussians for rendering, where colors are composited using alpha blending in a front-to-back order. Compared to NeRF-based methods  [5, 26], Gaussian Splatting offers higher reconstruction fidelity, faster rendering, and lower memory consumption during training. The rendering process can be formulated as $r=g(\theta,c)$, where $\theta$ denotes the 3D Gaussians, $g$ represents the renderer, $c$ captures the camera parameters, and $r$ is the resulting image.
Score Distillation Sampling (SDS)  [32] optimizes a 3D representation $\theta$ using a pre-trained 2D diffusion model $\phi$. The objective is to refine the 3D representation so that images rendered from various camera projections closely match those generated by the pre-trained 2D text-to-image diffusion model. The optimization process is guided by the SDS gradient, formulated as:
$\nabla_{\theta}\mathcal{L}_{\text{SDS}}(z;y)=\mathbb{E}_{t,\epsilon}\Big{[}\omega(t)\big{(}\hat{\epsilon}_{\phi}(z_{t},t,y)-\epsilon\big{)}\frac{\partial z}{\partial\theta}\Big{]},$ (1)
where, $\hat{\epsilon}_{\phi}$ represents the noise predicted by the model $\phi$, and $z_{t}$, $t$, and $y$ correspond to the noisy image, timestep, and text embedding, respectively. $\epsilon$ denotes Gaussian noise, and $\omega(t)$ is the weighting function applied at each timestep. SDS treats $\hat{\epsilon}_{\phi}(.)$ as a score function that estimates the sampled noise $\epsilon$ present in $z_{t}$ at a specific noise level $t$, conditioned on the text embedding $y$. The optimization minimizes the discrepancy between the predicted noise and the actual noise, guiding the 3D representation $\theta$ to produce renderings that align with the given text prompt.
Figure 3: Overall pipeline of DecompDreamer. Given a text prompt, a VLM generates an image and a scene graph. The scene graph identifies objects and their relationships, while the image provides spatial structure. Using this information, we perform coarse initialization followed by a progressive optimization strategy. Initially, optimization focuses on joint relationship modeling before gradually shifting towards targeted and view-aware object optimization, resulting in high-fidelity 3D objects with strong object disentanglement.

### v2

# DecompDreamer: A Composition-Aware Curriculum for Structured 3D Asset Generation

@@ Abstract
Current text-to-3D methods excel at generating single objects but falter on compositional prompts. We argue this failure is fundamental to their optimization schedules, as simultaneous or iterative heuristics predictably collapse under a combinatorial explosion of conflicting gradients, leading to entangled geometry or catastrophic divergence. In this paper, we reframe the core challenge of compositional generation as one of optimization scheduling. We introduce DecompDreamer, a framework built on a novel staged optimization strategy that functions as an implicit curriculum. Our method first establishes a coherent structural scaffold by prioritizing inter-object relationships before shifting to the high-fidelity refinement of individual components. This temporal decoupling of competing objectives provides a robust solution to gradient conflict. Qualitative and quantitative evaluations on diverse compositional prompts demonstrate that DecompDreamer outperforms state-of-the-art methods in fidelity, disentanglement, and spatial coherence.
Figure 1: Illustration of high-quality compositional 3D assets generated by DecompDreamer for complex text prompts. Existing methods often miss objects or distort spatial relationships, while DecompDreamer accurately captures geometry and preserves inter-object layout.
@@ 1 Introduction
The generation of complex, multi-object 3D scenes from natural language Bai et al. (2023); Gao et al. (2024); Ge et al. (2024); Po & Wetzstein (2024), or compositional text-to-3D, is fundamentally a high-dimensional combinatorial optimization problem. The task requires discovering a single plausible configuration from a virtually infinite search space of object geometries, attributes, and spatial relationships, guided only by the sparse constraints of a text prompt. The complexity of this search space grows exponentially with the number of objects and relationships, rendering naïve optimization strategies intractable. Common failure modes, such as misattributing colors to the wrong objects or generating physically implausible spatial arrangements, are the predictable outcomes of navigating this extraordinarily complex and ill-posed optimization landscape.
To make this problem tractable, existing methods Ge et al. (2024); Gao et al. (2024) rely on various optimization heuristics; however, they are often limited by a shared underlying pathology: conflicting gradients Yu et al. (2020). Compositional generation can be viewed as an implicit multi-task learning Caruana (1997); Kendall et al. (2018) problem where each object and relationship constitutes a distinct “task”. When multiple objectives, such as generating a high-fidelity object, binding its attributes correctly, and positioning it relative to others, are pursued simultaneously, the resulting supervisory signals can interfere destructively. Holistic methods Poole et al. (2023); Liang et al. (2024); Xiang et al. (2024) that treat the scene as a single entity suffer from severe gradient conflicts between implicit object tasks. More advanced methods Gao et al. (2024); Zhou et al. (2024) that decompose the scene using spatial layouts or semantic scene graphs mitigate this to some degree, but their reliance on simultaneous or joint optimization schedules re-introduces the problem. They attempt to solve for global structure and local detail concurrently, leading to an intractable multi-task problem that struggles to scale and often results in entangled or blended geometry.
This analysis reveals that the critical, unaddressed bottleneck in compositional generation is not just the static decomposition of the scene, but the dynamic scheduling of the optimization process. In this paper, we propose DecompDreamer, a framework that introduces a staged, composition-aware optimization strategy that resolves this fundamental challenge. Our approach is framed as an implicit curriculum Bengio et al. (2009) that transforms the intractable problem of joint generation into a sequence of manageable sub-problems. Inspired by the traditional 3D art workflows, we establish a scene’s global structure before refining individual details Chopine (2011), our method operationalizes this into a two-stage optimization schedule. In Stage One: Joint Relationship Modeling, we first prioritize the establishment of a coherent structural scaffold by focusing the optimization on inter-object relationships. This solves for the low-frequency components of the scene, effectively constraining the search space for the next stage. In Stage Two: Targeted and Disentangled Refinement, with the global structure fixed, we shift focus to refining the high-frequency details of individual objects. This stage employs targeted relational losses and view-aware supervision to enhance fidelity and enforce disentanglement without disrupting the previously established coherence.
By temporally decoupling the optimization of competing objectives, our staged approach provides a robust, structural mechanism for mitigating gradient conflict. This allows DecompDreamer to generate high-fidelity, well-disentangled 3D assets from highly complex prompts involving numerous objects and intricate interactions.
@@
  • We introduce a novel staged optimization strategy, framed as an implicit curriculum, that transforms the intractable problem of joint compositional generation into a sequence of manageable sub-problems. This “Composition-Aware” schedule first establishes global relational structure and then refines local object detail, providing a robust solution to the gradient conflicts that limit prior work.
  • We propose a set of techniques for the detail-refinement stage, including targeted relational optimization and view-aware supervision, that achieve strong object disentanglement by preventing feature blending and ensuring geometric consistency without disrupting the globally coherent structure established in the first stage.
  • We demonstrate through extensive qualitative and quantitative comparisons that our staged optimization paradigm significantly outperforms state-of-the-art methods in fidelity, coherence, and scalability, particularly on complex prompts with numerous objects and intricate interactions.
2 Theoretical Foundations
Figure 2: A visual taxonomy of optimization heuristics. (a) Holistic uses a single global loss. (b) Simultaneous applies all losses concurrently. (c) Iterative applies losses in a sequential loop. (d) Our staged curriculum temporally decouples relational and object losses
2.1 The Combinatorial Complexity of the Task
Compositional text-to-3D generation is a high-dimensional combinatorial optimization problem Du et al. (2020): a search for an optimal scene configuration $S$ that maximizes a plausibility function $P(S|T)$ given a text prompt $T$. The configuration $S$ is a complex tuple of objects $\{O_{i}\}$, attribute bindings $A:\{O_{i}\rightarrow\{a_{ij}\}\}$, and relational predicates $R:\{(O_{i},O_{j})\rightarrow\{r_{k}\}\}$. The search space for $S$ is combinatorially vast. For a prompt like “A blue bird sitting on a black branch of a white tree,” the model must solve the discrete assignment problem of binding attributes to entities. An incorrect binding can still yield a plausible rendering from certain views, creating a poor local minimum. Furthermore, satisfying relational predicates like “sitting on” imposes strong geometric constraints on the joint parameter space of the objects. The number of such interdependent variables grows exponentially, making the problem intractable without effective heuristics to structure and prune the search.
2.2 A Taxonomy of Optimization Heuristics
The intractability of this problem necessitates the use of optimization heuristics. We propose a taxonomy of prior work based on how each strategy applies object-level ($\mathcal{L}_{\text{obj}}$), relational ($\mathcal{L}_{\text{edge}}$), and scene-level ($\mathcal{L}_{\text{scene}}$) objectives.
Heuristic 1: Holistic Optimization. This baseline approach Xiang et al. (2024); Liang et al. (2024); Nath et al. (2024); Lin et al. (2023); Shi et al. (2024); Zhu et al. (2023); Chen et al. (2023) treats the scene as a monolithic entity, driven by a single objective, $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{scene}}$. It imposes no explicit compositional structure and is highly susceptible to conflicting gradients from the multiple implicit tasks, leading to entangled representations and incorrect attribute binding.
Heuristic 2: Simultaneous Decomposed Optimization. This heuristic decomposes the scene but optimizes all components concurrently via a weighted sum: $\mathcal{L}_{\text{total}}=w_{1}\mathcal{L}_{\text{obj}}+w_{2}\mathcal{L}_{\text{edge}}+w_{3}\mathcal{L}_{\text{scene}}$. Methods like Gala3D use strong structural priors (e.g., 3D bounding boxes) that prevent catastrophic failure. However, the simultaneous optimization of object geometry and scene layout, combined with the rigidity of the priors lead to a predictable limitation: stable but suboptimal convergence. This is particularly evident in complex interactions that cannot be fully captured within simple spatial containers, such as overlapping or interdependent object relationships (e.g., “knight riding horse”).
Heuristic 3: Iterative Decomposed Optimization. This approach also decomposes the scene but applies objectives in an alternating loop within each iteration (e.g., (i) $\mathcal{L}_{\text{obj}}$, then (ii) $\mathcal{L}_{\text{edge}}$, etc.). This heuristic, used by methods like GraphDreamer Gao et al. (2024), DreamHOI Zhu et al. (2024) and CompGS Ge et al. (2024), is vulnerable to optimization divergence at scale. As scene complexity grows, the combinatorial explosion of conflicting gradients creates a chaotic “tug-of-war,” causing the averaged gradient to become nonsensical. This leads the process to diverge rather than converge, resulting in a total failure of generation.
Heuristic 4: Staged Decomposed Optimization (DecompDreamer). This paradigm, which we introduce, is a structured, multi-stage curriculum Bengio et al. (2009); Caruana (1997); Kendall et al. (2018) that avoids optimizing the entire scene at once. Instead, it first tackles coherent sub-problems by prioritizing relational objectives ($\mathcal{L}_{\text{joint}}$) to establish a global scaffold (Stage 1: Structure-focused). Only then does it shift focus to refining individual components with object-level and targeted losses ($\mathcal{L}_{\text{obj}}$, $\mathcal{L}_{\text{target}}$) that preserve this structure (Stage 2: Detail-focused). This “structure-then-detail” approach ensures that the gradients within each stage are more self-consistent and minimally conflicting. By temporally decoupling competing objectives, it avoids the optimization divergence plaguing iterative methods, ensuring stable and scalable convergence. Our ablation study (Figures 5, 7) provides empirical validation, showing a non-staged variant fails to produce structurally consistent results.
@@ 3 Related Work
3.1 Text-to-3D Generation
Methods for generating 3D content from text descriptions are broadly categorized into two paradigms: 3D supervised (feed-forward) Cheng et al. (2023); Jun & Nichol (2023); Nichol et al. (2022); Wei et al. (2023); Xiang et al. (2024)techniques and 2D lifting (optimization-based) techniques Chen et al. (2024; 2023); Huang et al. (2023); Poole et al. (2023); Wang et al. (2023); Zhu et al. (2023). 3D supervised methods train models directly on paired text and 3D data, enabling the generation of high-quality 3D assets. However, the efficacy of these approaches is fundamentally limited by the availability and diversity of large-scale 3D datasets Deitke et al. (2023). As most existing datasets consist primarily of single-object models, these methods struggle to generalize to complex, compositional scenes involving multiple objects and interactions. 2D lifting techniques circumvent the need for large 3D datasets by instead leveraging powerful, pre-trained 2D diffusion models to guide the optimization of a 3D representation Mildenhall et al. (2021); Kerbl et al. (2023). The seminal work in this area, DreamFusion Poole et al. (2023), introduced Score Distillation Sampling (SDS), a method for aligning rendered views of a 3D model with a 2D diffusion prior. This optimization-based paradigm inspired numerous advancements Liang et al. (2024); Lin et al. (2023); Tang et al. (2023a; b); Wang et al. (2023). However, these foundational methods typically treat the entire scene as a single, monolithic entity. While effective for individual objects, they lack explicit mechanisms for modeling inter-object relationships and thus underperform significantly on compositional prompts.
3.2 Compositional 3D Generation
To address the limitations of monolithic generation, a dedicated subfield Dhamo et al. (2021); Bai et al. (2023); Po & Wetzstein (2024); Zhai et al. (2024) has emerged focusing on compositional scenes, which can be analyzed using the framework we introduced in Section 2.2. An initial wave of research exemplifies what we classify as Simultaneous Decomposed Optimization (Heuristic 2). These methods integrate explicit layout priors, such as non-overlapping 3D bounding boxes, to provide object-level guidance. Gala3D Zhou et al. (2024), for instance, is a canonical example that adopts a layout-guided approach. While these methods effectively prevent catastrophic failures, their reliance on rigid spatial representations often results in stable but suboptimal convergence. Empirical results indicate that they struggle to capture complex interactions where object geometries must overlap plausibly (e.g., “a knight riding a horse”). Subsequent work moved towards more flexible semantic decomposition strategies, often employing Iterative Decomposed Optimization (Heuristic 3). GraphDreamer was a key advance, proposing the use of a scene graph to model objects and their relationships separately. However, its iterative optimization schedule is susceptible to optimization divergence: as the number of objects increases, the combinatorial explosion of conflicting gradients can lead to catastrophic failures. This is consistent with empirical observations, where GraphDreamer often fails on prompts containing more than four objects, producing incoherent or nonsensical outputs. Our work, DecompDreamer, introduces a novel approach with Staged Decomposed Optimization (Heuristic 4). Like recent methods Gao et al. (2024); Zhou et al. (2024), we use 3D Gaussians and a semantic scene graph. However, we address the optimization-based failures of prior heuristics by introducing a principled curriculum that temporally decouples competing objectives. This staged approach, which shifts from modeling inter-object relationships to refining disentangled objects, is designed to ensure stable and scalable convergence where other heuristics fail.
@@ 4 Preliminaries
Gaussian Splatting Kerbl et al. (2023) represents a scene as a set of 3D Gaussians defined by means, covariances, colors, and opacities. These are projected as 2D Gaussians and composited via alpha blending in front-to-back order. Compared to NeRF-based methods Barron et al. (2022); Müller et al. (2022), this approach offers higher fidelity, faster rendering, and lower memory usage. The rendering process is defined as $r=g(\theta,c)$, where $\theta$ is the set of Gaussians, $g$ is the renderer, and $c$ the camera parameters.
Score Distillation Sampling (SDS) Poole et al. (2023) optimizes a 3D representation $\theta$ using guidance from a pre-trained 2D diffusion model $\phi$. At each iteration, an image $z$ is rendered from $\theta$ using a random camera viewpoint and a view-dependent text prompt $y_{\psi}$, where $\psi$ denotes the sampled azimuth angle. The diffusion model provides a supervisory gradient to update $\theta$ such that its renderings align with the prompt. The optimization is guided by the SDS gradient:
$\nabla_{\theta}\mathcal{L}_{\text{SDS}}(z;y_{\psi})=\mathbb{E}_{t,\epsilon}\left[\omega(t)\left(\hat{\epsilon}_{\phi}(z_{t},t,y_{\psi})-\epsilon\right)\frac{\partial z}{\partial\theta}\right],$ (1)
where $\hat{\epsilon}_{\phi}$ is the noise predicted by the diffusion model, $z_{t}$ is the noisy version of the rendered image, and $\omega(t)$ is a weighting function. This process distills the knowledge of the 2D model into the 3D representation. In our work, we replace the standard SDS loss with supervision from a flow model AI (2023), which learns a direct vector field from noisy to clean images for improved gradient stability.
Figure 3: Overview of the DecompDreamer pipeline. Given a text prompt, a VLM generates a scene graph to guide a coarse initialization. The core of our method is a composition-aware optimization curriculum that first models joint relationships to build a coherent structure, then refines individual objects to produce high-fidelity, disentangled 3D assets.